OpenStudy (megannicole51):

Calculate the integral e^(2x)sin(2x)dx

OpenStudy (anonymous):

Is sin(2x) also raised to the power?

OpenStudy (anonymous):

is this one of those parts in a circle problems?

OpenStudy (megannicole51):

nope just e nope its just studying for my calc 2 exam and im not sure if i can just solve this any way...such as integration by parts?

zepdrix (zepdrix):

@satellite73 yes :3

OpenStudy (anonymous):

parts and then i think you go around in circles i.e. get back what you started with, then divide by 2

OpenStudy (anonymous):

try integrating by parts and see what happens i loathe integration with a passion, mostly because it is a complete show off waste of time

OpenStudy (megannicole51):

how can i solve this integral quickly and efficiently

OpenStudy (megannicole51):

keep in mind I'm studying for exam:D

OpenStudy (anonymous):

actually i am not sure if it makes any difference what you choose for \(u\) and \(dv\) you can try \(u=e^x, du = e^xdx, dv = \sin(x), v=-\cos(x)\) and get \[-\cos(x)e^x+\int e^x\cos(x)dx\]

OpenStudy (anonymous):

no there is no real quick way then parts again

OpenStudy (megannicole51):

correct...and then solve for the integral of e^(x)cos(x)

OpenStudy (megannicole51):

can i just put up the questions and u can check my answers?

OpenStudy (anonymous):

that means i would have to do them and i would make a mistake quickest bet is to use wolfram to check

OpenStudy (megannicole51):

okay thanks

OpenStudy (anonymous):

but you get the idea right? when you use parts again, you are going to get the original integral back, then you solve for it

OpenStudy (megannicole51):

yeah

OpenStudy (anonymous):

\[\int e^x\sin(x)=-\cos(x)e^x+\int e^x\cos(x)dx=\]\[-\cos(x)e^x+e^x\sin(x)-\int e^x\sin(x)dx\]

OpenStudy (anonymous):

i dropped the \(2x\) so i guess you have to be more careful with the constants

OpenStudy (megannicole51):

where did u get the e^(x)sin(x) on the left side?

OpenStudy (anonymous):

it was the original question

OpenStudy (anonymous):

i mean the original integral

OpenStudy (megannicole51):

why would u add the original equation?

OpenStudy (anonymous):

you end up with \[\int e^x\sin(x)dx=-\cos(x)e^x+e^x\sin(x)-\int e^x\sin(x)dx\] and so \[2\int e^x\sin(x)=-\cos(x)e^x+e^x\sin(x)\] then divide by 2

OpenStudy (anonymous):

i.e. when you integrate by parts twice, the second time you get the minus the original integral back, then you solve for it

OpenStudy (megannicole51):

okay

OpenStudy (anonymous):

more or less clear?

OpenStudy (megannicole51):

nope but its fine ill ask someone tomorrow...

OpenStudy (anonymous):

we can do it slow if you like, this one is not so hard

OpenStudy (megannicole51):

im beyond frustrated with this work....they gave us a study sheet but no answer key.... no its fine thank you though

zepdrix (zepdrix):

aw :c

OpenStudy (anonymous):

ok yw but just to try \[\int e^x\sin(x)dx=-\cos(x)e^x+e^x\sin(x)-\int e^x\sin(x)dx\] is like \[Y=-\cos(x)e^x+e^x\sin(x)-Y\] solve for \(Y\)

OpenStudy (anonymous):

add \(Y\) to both sides, divide both sides by 2 that will get it

OpenStudy (megannicole51):

im guna work it out and see if i get anything close to the right answer...

OpenStudy (megannicole51):

I dont understand how youre explaining it im sorry

OpenStudy (megannicole51):

if I have the integral of 2cos(2x)*e^(2x) can I take out both 2s and put them outside the integral? @satellite73

zepdrix (zepdrix):

Yes,\[\Large \int\limits 2\cos(2x) e^{2x}\;dx \quad=\quad 2\int\limits \cos(2x)e^{2x}\;dx\] I'm not sure what you mean by `both` 2s. :o

OpenStudy (megannicole51):

you answered my question...thank you:)